# Inference-Time Text-to-Video Alignment with Diffusion Latent Beam Search

### v1

@@ Appendix C List of Prompts
Appendix D Detailed Analysis on Reward Function for Perceptual Video Quality
@@
Appendix E Prompt of AI Feedback
@@
Appendix F Calibrating Reward to Preference Feedback from GPT-4o
Figure 12 shows the two dimensional histogram and correlation between reward function and AI feedback from GPT-4o (OpenAI, 2023), and Figure 13 represent the coefficient of calibrated reward designed for GPT-4.
@@ Appendix F Calibrating Reward to Preference Feedback from GPT-4o
Figure 13: Coefficients of calibrated reward with GPT-4o.
Appendix G Qualitative Evaluation
We provide example videos among individual rewards and VLM calibrated rewards in Figure 14.
@@ Appendix G Qualitative Evaluation
Figure 14: We selected the video with the highest reward from 64 randomly generated videos. Videos chosen based on subject consistency or motion smoothness tend to lack motion, while those selected based on dynamic degree or imaging quality often fail to adhere to the given prompt. Additionally, evaluations based on dynamic degree may also lead to a loss of temporal consistency. Text-video consistency often exhibits a high correlation with VLM-based evaluation and, among individual metrics (Figure 3), is relatively effective in capturing the overall quality of a video. However, it may overlook certain aspects, such as frame-wise artifacts. Videos selected by VLM-calibrated rewards demonstrate a more balanced quality.
Appendix H Additional Results for Inference-Time Text-to-Video Alignment
@@
Figure 16 the performance of inference-time search on DEVIL-medium that includes the results with $KB=64$. We can observe that the increasing trends still continue.
@@
Appendix I Extended Related Works
Aligning Diffusion Models via Finetuning   Alignment by finetuning text-conditioned models has been investigated for image (Lee et al., 2023b) and video (Furuta et al., 2024b; Wu et al., 2024c) generation. Typically, LoRA (Hu et al., 2021) in a backbone model (Çiçek et al., 2016) is finetuned through policy gradient (Black et al., 2024; Fan et al., 2023; Zhang et al., 2024a), direct preference optimization (Yuan et al., 2024; Na et al., 2024; Yang et al., 2023; Liang et al., 2024; Wallace et al., 2023a), reward-weighted regression (Dong et al., 2023), or direct reward gradient (Prabhudesai et al., 2024; Clark et al., 2024; Wu et al., 2024c; Li et al., 2024a). Some train an extra model for better initial noise space (Qi et al., 2024; Ahn et al., 2024; Zhou et al., 2024). In contrast, we focus on the search over the denoising process at inference time, which does not require any model updates and may not degrade the original performance.
Evaluation of Text-to-Video Generation   While there are several conventional metrics for video generation (or the one repurposed from image generation) such as SSIM (Wang et al., 2004), IS (Salimans et al., 2016), LPIPS (Zhang et al., 2018), or FVD (Unterthiner et al., 2019), those are not always suitable to evaluate how the quality of contents in video is, which is much more emphasized in text-to-video generation (Wu et al., 2024a). It has been a long-standing challenge to comprehensively and semantically evaluate the dynamics of contents or physical commonsense in generated videos (Bansal et al., 2024; Liao et al., 2024). To deal with that, VBench (Huang et al., 2024b) is recently proposed as a suite of holistic evaluations for text-to-video generation to reflect the perceptual aspect of the quality, such as consistency, smoothness, aesthetics of contents or text–video alignment. Moreover, inspired by the success in LLMs (Bai et al., 2022; Lee et al., 2023a; Furuta et al., 2024a), we could leverage VLMs, which become more capable these days, as a proxy of human evaluation of the contents (Wu et al., 2024c); by finetuning CLIP-based models (He et al., 2024; Radford et al., 2021; Ma et al., 2022), or prompting GPT-4o (OpenAI, 2023) or Gemini (Gemini Team, 2023). Our paper adopts AI feedback from VLMs as an alternative to human rater, and proposes a recipe to calibrate a reward to other sources of feedback (such as AI or human feedback), by considering a linear combination of fine-grained metrics.
Appendix J Details of Sample Diversity
We compute the pairwise distances of ViCLIP (Xu et al., 2021) embeddings for video generation to quantify the diversity in videos, inspired by the approach for evaluating diversity in images (Kim et al., 2025). Specifically, given $N$ generated video samples, we first extract ViCLIP embeddings $\mathbf{v}^{\text{video},(i)}$ for each sample $i$. The pairwise diversity score is then computed as the mean pairwise distance:

### v2

@@ Appendix F List of Prompts
DEVIL-very-high
  1. 1.
    Classical style of a horse partaking in an ancient chariot race, scenes switching quickly from cheering crowds to close-ups of intense wheel clashes.
  2. 2.
    High-speed shots of a volcanic eruption engulfing a tropical island, with lava fountains spewing molten rock and the environment transforming from idyllic paradise to hellish landscape of ash and fire.
  3. 3.
    A thrilling scene of rural mountain biking extreme sports, starting from the early morning cycling adventure, transitioning to the intense chase through fields and forests, and ending with cheers and celebrations under the sunset.
  4. 4.
    Neon-lit streets pulse with energy as vehicles engage in a high-octane pursuit, transitioning seamlessly from chaos to calculated evasion. Against the backdrop of a setting sun, the chase intensifies, each turn a heartbeat away from capture.
  5. 5.
    A fighter jet dodging rapid anti-air gunfire, quick maneuvers, tracer rounds visible.
  6. 6.
    Bear escaping a collapsing cave, rocks tumbling, dust rising, ((masterpiece)), ((best quality)), 8k, high detailed, ultra-detailed, bear, ((dark rocky textures)), sprinting, (echoing rumble), sudden movement
  7. 7.
    A courier on a bike weaves through traffic at breakneck speed, narrowly avoiding cars and pedestrians in a rush to make deliveries on time.
  8. 8.
    A hummingbird rapidly darting between vibrant flowers in a lush garden, with quick cuts to various close-up shots showcasing its rapid wing movement and agility.
  9. 9.
    Venetian gondola chase scene, narrow canals, historic buildings, urgent escape, ((masterpiece)), ((best quality)), 8k, high detailed, ultra-detailed, gondola, ((twisting canals)), (ancient architecture), (urgent paddling), cinematic chase.
  10. 10.
    Futuristic sports cars racing on a vertical loop track against a sci-fi cityscape, cars defying gravity, ((speed trails)), (dizzying heights), (spectacular crashes), the thrill of cutting-edge technology.
  11. 11.
    Cat rapidly zigzagging across a rooftop, avoiding swooping birds under a stormy sky.
  12. 12.
    A sequence of a cow performing acrobatic stunts over a series of colorful, abstract platforms that morph shapes.
  13. 13.
    The dog bursts through a thicket, darting from a foggy forest to a steep hillside, rocks crumbling under its paws as it charges towards a roaring river below.
  14. 14.
    Thundering across a vast desert plain, the elephants race over dunes and dodge sandstorms, before swiftly traversing through a rocky canyon, bounding over boulders and leaping across narrow ravines.
  15. 15.
    A giraffe navigating a city during a robot uprising, with quick cuts showing chaotic battles, explosions, and futuristic technology in a high-stakes escape scenario.
  16. 16.
    A horse leading a wild stampede across a stormy beach with waves crashing, depicted with swift, sweeping camera moves, cinematic composition.
  17. 17.
    Intense motorcycle escape from a volcanic eruption, with transitions from lava-filled landscapes to ash-clouded skies.
  18. 18.
    A futuristic robot uprising, ((lasers firing)), metallic drones, explosions, debris, ((screaming civilians)), dystopian cityscape.
  19. 19.
    Sheeps engaging in a high-speed pursuit through a cyberpunk city, the scene rapidly transitioning between neon-lit streets, bustling marketplaces, and towering skyscrapers.
  20. 20.
    A pulse-pounding sequence of a train barreling through a treacherous storm, the scene transitioning between lightning-lit skies and torrents of rain to flooded tracks and collapsing bridges.
  21. 21.
    A truck rushing away from a treacherous mountain pass during a blizzard, with sudden avalanches and rockslides adding to the danger.
  22. 22.
    A zebra sprinting across the busy lanes of Times Square in New York City, with scene transitions occurring quickly as it moves from iconic billboards to bustling sidewalks filled with tourists.
MovieGen
  1. 1.
    A green monster made of plants walks through an airport.
  2. 2.
    A marble goes through a glass cup, breaking it into pieces.
  3. 3.
    A droplet of water falling onto a hot surface, instantly evaporating into a wisp of steam that swirls gracefully into the air.
  4. 4.
    An old man wearing a green dress and a sun hat taking a pleasant stroll in Johannesburg South Africa during a beautiful sunset.
  5. 5.
    A person on a hoverboard colliding with a wall, the board stopping abruptly.
  6. 6.
    A toy robot wearing blue jeans and a white t-shirt taking a pleasant stroll in Johannesburg South Africa during a winter storm.
  7. 7.
    In a marathon race, a female athlete gradually sprints ahead of the male athletes.
  8. 8.
    A teenager eating a slice of pizza, cheese stretching as they pull it away.
  9. 9.
    A man in a suit fights monsters.
  10. 10.
    A dog made of ice melts completely in a hot summer day.
  11. 11.
    A truck right alongside a flowing river, capturing the movement of the water and the surrounding forest.
  12. 12.
    A group of skateboarders perform tricks on ramps and rails at a skate park, showcasing their skills.
  13. 13.
    A hot air balloon descending back to the ground.
  14. 14.
    Chef chopping onions in the kitchen for the preparation of the dish.
  15. 15.
    Zoom in shot to the face of a young woman sitting on a bench in the middle of an empty school gym.
  16. 16.
    The couple runs hand in hand to release a sky lantern, then watches it drift upward into the night sky, carried by the wind with the stars shining above.
  17. 17.
    Aerial view shot of a cloaked figure elevating in the sky between skyscrapers.
  18. 18.
    A softball player sliding safely into second base.
  19. 19.
    A giraffe in a lifeguard outfit, sitting atop a high chair and watching over a crowded pool.
  20. 20.
    A speed skater accelerating during a short track race.
Appendix G Detailed Analysis on Reward Function for Perceptual Video Quality
@@
Appendix H Prompt of AI Feedback
@@
Appendix I Further Results for Calibrating Reward to Preference Feedback
I.1 Basic Prompts
Figure 12 and Figure 13 shows the two dimensional histogram and correlation between reward function and AI feedback from Gemini [14] and GPT-4o [13], and Figure 10 represent the coefficient of calibrated reward designed for GPT-4o. The relative weighting of the dynamic degree shifts in accordance with the prompt-specified dynamics, reflecting a pattern similar to that observed in the reward calibration approach with Gemini (DEVIL-high, DEVIL-medium, and DEVIL-static denote prompt sets characterized by high, medium, and low levels of dynamics, respectively). GPT-4o exhibits a stronger inclination toward dynamics than Gemini.
Figure 10: Coefficients of calibrated reward with GPT-4o.
I.2 Challenging Prompts
This section describes the reward calibration procedure and results for two challenging prompt sets, DEVIL-very-high and MovieGen, which were introduced to evaluate our method with larger T2V models. Following the methodology from for reward calibration with Latte (see Section 4), we generated 64 videos per prompt using Wan 2.1-1.3B [20]. Consistent with observations in Appendix I.1, using solely Text-Video Consistency is insufficient to fully capture AI feedback from Gemini (Figure 14). We choose the combination of weights $w_{i}$ to maximize correlation with Gemini’s evaluations. The coefficients of calibrated weights are shown in Figure 11.
Figure 11: Coefficients of calibrated reward with Gemini.
Figure 12: 2D-histogram and correlation between reward function and AI feedback from Gemini.
@@ I.2 Challenging Prompts
Figure 14: 2D-histogram and correlation between reward function and AI feedback from Gemini for challenging prompt sets, DEVIL-very-high and MovieGen.
Appendix J Correlation between VLM and Human Evaluation
As mentioned in the prior research [44, 45], VLMs such as Gemini and GPT-4o have been shown to exhibit a high correlation with human evaluations compared to other existing metrics. As an experiment, we measured the correlation between the AI feedback from these VLMs and human labels in the TVGE [72] dataset. As shown in Figure 15, Gemini achieved a correlation of 0.49, and GPT-4o achieved 0.51. Consequently, optimizing for these VLM rewards is a valid way to improve human-perceived quality, rather than merely “gaming” the metrics.
Figure 15: Correlation between VLM outputs and human labels in the TVGE dataset.
Appendix K Qualitative Evaluation of Calibrated Reward
@@
We provide example videos among individual rewards and VLM calibrated rewards in Figure 16. Videos selected solely on a single metric can over-optimize one aspect while neglecting others, whereas those chosen via VLM-calibrated rewards exhibit a more balanced quality. For instance, videos chosen solely based on temporal consistency (i.e., subject consistency and motion smoothness) or frame-by-frame quality (i.e., aesthetic quality, imaging quality) tend to lack dynamic movement, whereas those selected based on dynamic degree often lose temporal consistency. Evaluations relying on a single metric also fail to reflect the given prompt in some cases. Text-video consistency, which often exhibits a high correlation with VLM-based evaluation among individual metrics (Figure 3), is relatively effective in capturing the overall quality of a video. However, it may overlook certain aspects, such as frame-wise artifacts. In contrast, videos selected using VLM-calibrated rewards exhibit a more balanced overall quality.
Figure 16: We select the video with the highest reward out of 64 randomly generated candidates for each prompt, drawn from DEVIL-high, DEVIL-medium, DEVIL-static, and MSRVTT-test (arranged from top to bottom). Videos chosen using VLM-calibrated rewards achieve a more balanced quality compared to those relying on any single metric. For instance, when subject consistency, motion smoothness, or aesthetic quality serves as the sole selection criterion, the resulting videos often lack dynamic movement, whereas prioritizing dynamic degree can compromise temporal consistency. Moreover, single-metric evaluations may occasionally fail to align with the intended prompt.
Appendix L Further Results for Diffusion Latent Beam Search
L.1 Scaling Search Budget with GPT-4o Calibrated Reward
We measure the performance in terms of a combinational reward calibrated to GPT-4o (Figure 17). DLBS improves all the calibrated rewards the best as the search budget $KB$ increases (especially $KB=16,32$), while BoN and GS, in some cases, eventually slow down or saturate the performance. Notably, an LA estimator with a small search budget ($KB=8,T^{\prime}=6$) is comparable to or even outperforms DLBS ($KB=32$).
Figure 17: Inference-time search on reward calibrated to GPT-4o.
L.2 Scaling Search Budget to Larger Regimes
Figure 18 and Figure 19 show the performance of inference-time search on DEVIL-medium and MSRVTT-test that includes the results with $KB=64$. We can observe that the increasing trends still continue.
L.3 Full Results for Scaling Trend of DLBS
Figure 20 demonstrates the scaling trend of DLBS, proportional to the search budget, under various choices of $K$. The results show that there is an optimal balance between the number of latent $K$ and the number of beams $B$ under the same budget. For instance, as we increase the budget to $KB=16,32$, we have a peak around $K=4,8,16$, which is about 25–50% of the budget. This implies that balancing possession and exploration of diffusion latents in DLBS helps search for the best outputs robustly.
L.4 Further Analysis on Lookahead Estimator
Figure 21 (Left) demonstrates that increasing the number of reward estimation steps $T^{\prime}$ in the LA estimator leads to improved reward prediction performance for $\mathbf{z}_{t}$ during the denoising process. This finding suggests that extending the LA steps enables a more effective search based on accurate reward predictions, particularly in the early stages of the denoising. As shown in Figure 21 (Right), enlarging the look-ahead horizon increases the reward gain to $T^{\prime}=6$; beyond this point, e.g., $T^{\prime}=20$ offers no significant benefit while multiplying the calculation. Accordingly, we fix $T^{\prime}=6$ in the main experiments, as it captures nearly all the attainable gains at minimal cost.
Figure 21: (Left) Comparison of the reward estimation error for different LA steps $\mathbf{T^{\prime}}$. We evaluate the reward predicted by the LA estimator, which projects $\mathbf{z}_{t}$ to $\tilde{\mathbf{z}}_{0|\tilde{t}(0)}$ in $T^{\prime}$ steps (Algorithm 2) and computes $r^{\prime}(\tilde{\mathbf{z}}_{0|\tilde{t}(0)})$, against the actual reward obtained by projecting $\mathbf{z}_{t}$ to $\mathbf{z}_{0}$ in $t$ steps using a DDIM sampler ($\eta=0.0$) and evaluating $r^{\prime}(\mathbf{z}_{0})$. (Right) Impact of $T^{\prime}$ on search performance. Reward improves rapidly up to $T^{\prime}=6$ but saturates thereafter; using $T^{\prime}=20$ offers no measurable gain. These results show that a modest $T^{\prime}$ is sufficient in practice.
L.5 Ablation Study for Diffusion Steps
Scaling Diffusion Steps  Figure 22 (Left) shows the performance when increasing the number of denoising steps $T$. Since DDIM exhibits fast convergence [42], BoN with larger $T$ does not improve the reward much. DLBS improves the performance when scaling denoising steps to $T=200$ more than BoN, which implies that DLBS benefits from larger computation and the other axes. However, as Figure 6 (Left) indicates, these gains are smaller than those obtained by widening the beam budget $KB$ or leveraging the LA estimator.
Range of Diffusion Steps for Search  We investigate which range of diffusion steps DLBS should be applied to for effective search. Unlike Kim et al. [15], which apply GS only during the initial 5–10 steps, our results in Figure 22 (Right) show that applying DLBS throughout all steps leads to substantially better performance. This suggests that applying DLBS entirely is more effective than focusing on early stage.
Figure 22: (Left) Scaling the denoising steps $T$. (Right) Range of denoising steps $t\in[50,0]$ to apply search methods. While Kim et al. [15] applies GS in the first 5–10 steps, DLBS over the entire diffusion steps yields the largest improvement.
L.6 Ablation Study for $\eta$ in DDIM scheduler
@@
L.7 Alignment–Diversity Tradeoff
Alignment for diffusion models can steer desirable outputs, but it is said that the diversity of generated samples or the performance of original models often degrade [10, 37]. While inference-time search does not change or degrade the model itself, we here compare the diversity of samples among BoN, GS, and DLBS. We measure the sample diversity as the mean pairwise distance of ViCLIP [51] embeddings to quantify the diversity in videos, inspired by the approach for evaluating diversity in images [58]. Specifically, given $N$ generated video samples, we first extract ViCLIP embeddings $\mathbf{v}^{\text{video},(i)}$ for each sample $i$. The pairwise diversity score is then computed as the mean pairwise distance: